# Joint distribution of Hecke eigenforms on $\mathbbm{H}^{3}$

### v1

@@ 1.1. Value distribution conjecture
The value distribution of eigenfunctions in the semiclassical limit is one of the main problems in analytic number theory and quantum chaos. Berry [Ber77] suggested that eigenfunctions of the Laplace operator for chaotic systems are modeled by random waves: in particular, eigenfunctions on a compact hyperbolic surface should have a Gaussian value distribution when the frequency grows, and the moments of an $L^{2}$-normalized eigenfunction should be given by the Gaussian moments. Such a behavior is analogously expected from automorphic forms on more general locally symmetric spaces, see [Hum18]. This conjectural behavior, called the Gaussian Moment conjecture, encapsulates the Quantum Unique Ergodicity conjecture.
This article considers the joint value distribution of Hecke-Maaß cusp forms on the hyperbolic 3-space $\mathbb{H}^{3}$, analogously to the recent results of Huang [Hua24, HHL24] for the hyperbolic plane $\mathbb{H}^{2}$. In the case of the hyperbolic $3$-space, we have the following conjecture (analogue of [Hua24, Conjecture 1.6]):
@@
where $\psi$ is a smooth compactly supported function. The following theorem is the main result of this paper and a smoothed analogue of the joint value equidistribution conjecture; it is analogous to [HHL24, Theorem 1.4] in the case of the Poincaré upper-half plane $\mathbb{H}^{2}$.
@@ 2.1. Hyperbolic 3-space and automorphic forms
We recall the theory of the hyperbolic $3$-space, for which a good reference is [EGM98]. Let
@@
The theory of Eisenstein series for the hyperbolic $3$-space can be found in [EGM98] or [Hei90]. In order to write the Fourier expansion of the Eisenstein series let $\zeta_{K}(s)$ denote the Dedekind zeta function for $K$ and
@@
as proven in e.g. [Hei90, p. 102]. Here $\sigma_{s}(\omega)$ denotes the generalized divisor function
@@
and $\langle\mu,z\rangle=\Re(\overline{\mu}z)$. A different way to define the Eisenstein series is
and then $E(P,s)=\frac{|\mathcal{O}_{K}^{*}|}{2}E_{\infty}(P,s)$, see [EGM98, p. 232]. In the three-dimensional case the Hecke operators are defined as follows: if $n\in\mathcal{O}_{K}\setminus\{0\}$ we define $\mathcal{M}_{n}$ to be the set of all matrices of the form $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)$, adbc=nad-bc=n. Then for $f$ being a $\Gamma$-invariant function the Hecke operator $T_{n}$ is given by
@@
with the inverses being taken in the skew field of quaternions. The theory for Hecke operators for Bianchi groups is developed in [Hei90]. However, in contrast to Heitkamp we have incorporated the factor $1/\sqrt{N(n)}$ in the definition of the Hecke operator.
The Fourier expansion of a cusp form is given in [Hei90, Lemma 16.1]. Let $f$ be an automorphic form with eigenvalue $\lambda=1-s^{2}$. Then we have the Fourier expansion at the cusp $\infty$ given by
@@
which should be compared to [Hei90, Korollar 17.6], noting that the convention therein is the arithmetic normalization $\rho_{f}(1)=1$ and therefore a non-$L^{2}$-normalized form. We emphasize that the numerator $32\pi^{4}$ in [Hei90, p. 125, last equation] should instead be $128\pi^{4}$.
@@ 2.2. Preliminary results
A central tool in our method is the spectral expansion and the Parseval equality, which reads as follows [EGM98, Section 6.3, Theorem 3.4].
@@
This is in line with [CFR21, (6.5)], with the precise computation of the “constants" appearing therein. See also the very recent paper [Nel25] explaining bounds on triple products in the real case of the quantum unique ergodicity conjecture.
For each $i\in\{1,2,3\}$, let $\pi_{i}$ be a cuspidal spherical representation in which $\phi_{i}\in\pi_{i}$ is a spherical vector. By the Ichino formula [Mar14, Theorem 4], there is an absolute constant $C$ such that
@@
We embed $\pi_{1}$ and $\pi_{2}$ in their respective Whittaker models: denote by $\mathcal{W}(\pi_{i},\psi)$ the Whittaker model of $\pi$ with respect to the character $\psi$ [Bum97, Section II.2.8]. Let $W_{i}\in\mathcal{W}(\pi_{i},\psi)$ be a Whittaker function corresponding to $\phi_{i}$. We embed $\pi_{3}$ in its induced model $\mathcal{I}_{3}$ as explained in [Bum97, Theorem II.2.7.1], and denote $f_{3}\in\mathcal{I}_{3}$ be the vector corresponding to $\phi_{3}$, which is given by [Bum97, (5.13) and (5.22)], viz.
@@
for all $g=n(x)a(y)k$ decomposed according to the Iwasawa decomposition of $G$, where $n(x)$ is the unipotent matrix with upper-right entry $x$ and $a(y)$ is the diagonal matrix parametrized by $y$. Here, $\lambda=s(1-s)$ relates the spectral parameter $s$ to the eigenvalue $\lambda$, and $s=\tfrac{1}{2}(s_{1}-s_{2}+1)=1+it$. By using [MV10, Lemma 3.4.2] also stated in [Mar14, Proposition 7], we can express in these terms the sought integral:
@@
We now follow the strategy used in [Mar14, Proposition 6] to deal with the integral over $K$. We examine the action of the central torus $M=K\cap Z$ on $W_{i}$ and $f_{3}$, for which these are eigenfunctions. Concerning $W_{i}$, the left action under $K$ is given in [Bum97, Proposition II.2.8.1] (it has a weight $k_{i}$ in the notations therein), and the right action under $Z$ is given in [Bum97, p. 244] for diagonal matrices (it has weight $\mu_{i}$); therefore the action under $M=K\cap Z$ has necessarily weight $k_{i}=\mu_{i}$. As for $f_{3}$, it has a given weight on both sides by the explicit computations from [Bum97, (5.22)].
@@
We appeal to the explicit knowledge of the Whittaker functions in our case [Mar14, Proposition 8, with weights all zero and therefore $m=0$], which is also given in [Mar14, (22), in the case $k=0$]. We have
@@
We input these values in $\mathcal{T}$ and use the inversion formula given in [GR07, 6.576] to obtain
@@
For every $t\in\mathbb{R}$, denote $E_{t}:=E(\cdot,it)$. For every $t,s\in\mathbb{R}$ and every cusp forms $f$ and $g$, we have
@@
We start by $\langle u_{k}g,E_{t}\rangle$, which we will unfold by inputting the Fourier expansion of each cusp form and Eisenstein series. This generalizes [Hei90, Section 17.2], where the computation is restricted to the case $u_{k}=g$. By using the Fourier expansions (11) of $u_{k}$ and $g$, we have
@@
By executing the integration of $e(\langle\tilde{\mu}-\tilde{\nu},z\rangle)$ over $z$ first, we select only the case where $\mu=\nu$ with a weight $\mathrm{vol}(P_{\mathfrak{o}})$; indeed by [Hei90, p. 122] we have
@@
By executing the integration of $e(\langle\tilde{\mu}-\tilde{\nu},z\rangle)$ over $z$ first and using (28), we only pick up the terms where $\mu=m$ with a weight $\mathrm{vol}(P_{\mathfrak{o}})$, and this in particular removes the constant terms of the Eisenstein series’ expansion. We are therefore left with
@@
$\langle E_{t}f,E_{s}\rangle$ $=\rho_{f}(1)2^{2-(1-it)}\frac{\sqrt{|d_{K}|}}{2}\frac{|O^{\times}|^{2}}{\xi_{K}(it)\xi_{K}(is)}\left(\frac{4\pi}{\sqrt{|d_{K}|}}\right)^{-it}$ (32)
$\quad\times\frac{L(\frac{it-is+1}{2},u_{j})L(\frac{it+is+1}{2},u_{j})}{\Gamma(it)}\prod_{\pm,\pm}\Gamma\left(\frac{it\pm(is-1)\pm it_{j}-1}{2}\right)$ (33)
@@
Assing [Ass19a] recently proved a stronger bound on sup-norm for Eisenstein series with exponent $3/8$, matching the quality of [You18]; however the above lemma is sufficient to our purposes and we provide an elementary proof for completeness.
@@
We follow closely [You18]. Consider the Fourier expansion given in (10) and the relation $E=\frac{|\mathcal{O}_{K}^{*}|}{2}E_{\infty}$. After applying the Cauchy-Schwarz inequality, it suffices to bound
@@
To simplify the notation, denote $L=4\pi r/\sqrt{|d_{K}|}$ and notice that both $L$ and $1/L$ are bounded in terms of $\Omega$. Applying the uniform asymptotic expansion due to Balogh [Bal67] for the $K$-Bessel function, we obtain the estimates
$\cosh(\pi t/2)K_{it}(u)\ll\begin{cases}t^{-1/4}(t-u)^{-1/4},\quad&\text{if }0<u<t-Ct^{1/3},\\ t^{-1/3},\quad&\text{if }|u-t|\leq Ct^{1/3},\\ u^{-1/4}(u-t)^{-1/4}\exp\Big{(}-c(\frac{u}{t})^{3/2}\big{(}\frac{u-t}{t^{1/3}}\big{)}^{3/2}\Big{)},\quad&\text{if }u>t+Ct^{1/3}.\end{cases}$ (36)
@@
using Shiu’s bound [Shi80] on the function $\tilde{\sigma}_{0}(n)=\frac{1}{\sqrt{n}}\sum_{|\omega|^{2}=n}\sigma_{0}(\omega)$. These terms then give the same bound as (37), plus an additional term that is $\ll_{\Omega}t^{-1/3+\varepsilon}$. The contribution from the remaining range with $u>2t$ is easily bounded thanks to the exponential contribution in (36). ∎
@@ 3. Proof of the theorem
We have now all the tools necessary to prove Theorem 1.2. We follows the strategy of [HHL24].
@@ 3.1. Spectral expansion and first truncation
Using the Cauchy-Schwarz inequality, the fact that $\|\Delta^{\ell}\psi\|_{1}\ll 1$ and $\|E_{t}\|_{\infty}\ll t^{1+\varepsilon}$ from Lemma 2.7, we obtain $\langle\psi,u_{k}\rangle\ll\lambda_{k}^{-\ell}$ for any $\ell>0$.
@@
$\langle\psi,f^{2}g\rangle$ $=\frac{\langle\psi,1\rangle\langle 1,f^{2}g\rangle}{\mathrm{vol}(\Gamma\backslash\mathbb{H}^{3})}+\delta_{t_{g}\ll t_{f}^{o(1)}}\langle\psi,g\rangle\langle 1,f^{2}g^{2}\rangle+\sum_{\begin{subarray}{c}u_{k}\neq g\\ t_{k}\ll\max(t_{f},t_{g})^{o(1)}\end{subarray}}\langle\psi,u_{k}\rangle\langle u_{k},f^{2}g\rangle$ (44)
@@ 3.2. Reduction to triple products
Many terms in (44) display inner products involving four forms. In order to make them amenable to the Watson-Ichino formula, we apply further the spectral expansion from Lemma 2.2 to the terms $u_{k}g$ and $E_{t}g$, which allows to reduce the expression to products of three forms at the cost of an extra spectral summation. We obtain

### v2

@@ 1.1. Value distribution conjecture
The value distribution of eigenfunctions in the semiclassical limit is one of the main problems in analytic number theory and quantum chaos. Berry [berry_regular_1977] suggested that eigenfunctions of the Laplace operator for chaotic systems are modeled by random waves: in particular, eigenfunctions on a compact hyperbolic surface should have a Gaussian value distribution when the frequency grows, and the moments of an $L^{2}$-normalized eigenfunction should be given by the Gaussian moments. Such a behavior is analogously expected from automorphic forms on more general locally symmetric spaces, see [humphries_equidistribution_2018]. This conjectural behavior, called the Gaussian Moment conjecture, encapsulates the Quantum Unique Ergodicity conjecture.
This article considers the joint value distribution of Hecke-Maaß cusp forms on the hyperbolic 3-space $\mathbb{H}^{3}$, analogously to the recent results of Hua, Huang and Li [huang_joint_2024, hua_joint_2024] for the hyperbolic plane $\mathbb{H}^{2}$. In the case of the hyperbolic $3$-space, we have the following conjecture (analogue of [huang_joint_2024, Conjecture 1.6]):
@@
where $\psi$ is a smooth compactly supported function. The following theorem is the main result of this paper and a smoothed analogue of the joint value equidistribution conjecture; it is analogous to [hua_joint_2024, Theorem 1.4] in the case of the Poincaré upper-half plane $\mathbb{H}^{2}$.
@@ 2.1. Hyperbolic 3-space and automorphic forms
We recall the theory of the hyperbolic $3$-space, for which a good reference is [emg]. Let
@@
The theory of Eisenstein series for the hyperbolic $3$-space can be found in [emg] or [hei]. A different way to define the Eisenstein series is by
and then $E(P,s)=\frac{|\mathcal{O}_{K}^{*}|}{2}E_{\infty}(P,s)$, see [emg, p. 232]. In order to write the Fourier expansion of the Eisenstein series let $\zeta_{K}(s)$ denote the Dedekind zeta function for $K$ and
@@
as proven in e.g. [hei, p. 102]. Here $\sigma_{s}(\omega)$ denotes the generalized divisor function
@@
and $\langle\mu,z\rangle=\Re(\overline{\mu}z)$. In the three-dimensional case the Hecke operators are defined as follows: if $n\in\mathcal{O}_{K}\setminus\{0\}$ we define $\mathcal{M}_{n}$ to be the set of all matrices of the form $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)$, adbc=nad-bc=n. Then for $f$ being a $\Gamma$-invariant function the Hecke operator $T_{n}$ is given by
@@
with the inverses being taken in the skew field of quaternions. The theory for Hecke operators for Bianchi groups is developed in [hei]. However, in contrast to Heitkamp we have incorporated the factor $1/\sqrt{N(n)}$ in the definition of the Hecke operator.
The Fourier expansion of a cusp form is given in [hei, Lemma 16.1]. Let $f$ be an automorphic form with eigenvalue $\lambda=1-s^{2}$. Then we have the Fourier expansion at the cusp $\infty$ given by
@@
which should be compared to [hei, Korollar 17.6], noting that the convention therein is the arithmetic normalization $\rho_{f}(1)=1$ and therefore a non-$L^{2}$-normalized form. We emphasize that the numerator $32\pi^{4}$ in [hei, p. 125, last equation] should instead be $128\pi^{4}$.
@@ 2.2. Preliminary results
A central tool in our method is the spectral expansion and the Parseval equality, which reads as follows [emg, Section 6.3, Theorem 3.4].
@@
This is in line with [nicole, (6.5)], with the precise computation of the “constants" appearing therein. See also the very recent paper [nelson_subconvexity_2025] explaining bounds on triple products in the real case of the quantum unique ergodicity conjecture.
For each $i\in\{1,2,3\}$, let $\pi_{i}$ be a cuspidal spherical representation in which $\phi_{i}\in\pi_{i}$ is a spherical vector. By the Ichino formula [marshall_triple_2014, Theorem 4], there is an absolute constant $C$ such that
@@
We embed $\pi_{1}$ and $\pi_{2}$ in their respective Whittaker models: denote by $\mathcal{W}(\pi_{i},\psi)$ the Whittaker model of $\pi$ with respect to the character $\psi$ [bump, Section II.2.8]. Let $W_{i}\in\mathcal{W}(\pi_{i},\psi)$ be a Whittaker function corresponding to $\phi_{i}$. We embed $\pi_{3}$ in its induced model $\mathcal{I}_{3}$ as explained in [bump, Theorem II.2.7.1], and denote $f_{3}\in\mathcal{I}_{3}$ the vector corresponding to $\phi_{3}$, which is given by [bump, (5.13) and (5.22)], viz.
@@
for all $g=n(x)a(y)k$ decomposed according to the Iwasawa decomposition of $G$, where $n(x)$ is the unipotent matrix with upper-right entry $x$ and $a(y)$ is the diagonal matrix parametrized by $y$. Here, $\lambda=\sigma(1-\sigma)$ relates the spectral parameter $\sigma$ to the eigenvalue $\lambda$, where $\sigma=\tfrac{1}{2}(s_{1}-s_{2}+1)=1+it$. By using [michel_subconvexity_2010, Lemma 3.4.2] also stated in [marshall_triple_2014, Proposition 7], we can express in these terms the sought integral:
@@
We now follow the strategy used in [marshall_triple_2014, Proposition 6] to deal with the integral over $K$. We examine the action of the central torus $M=K\cap Z$ on $W_{i}$ and $f_{3}$, for which these are eigenfunctions. Concerning $W_{i}$, the left action under $K$ is given in [bump, Proposition II.2.8.1] (it has a weight $k_{i}$ in the notations therein), and the right action under $Z$ is given in [bump, p. 244] for diagonal matrices (it has weight $\mu_{i}$); therefore the action under $M=K\cap Z$ has necessarily weight $k_{i}=\mu_{i}$. As for $f_{3}$, it has a given weight on both sides by the explicit computations from [bump, (5.22)].
@@
We appeal to the explicit knowledge of the Whittaker functions in our case [marshall_triple_2014, Proposition 8, with weights all zero and therefore $m=0$], which is also given in [marshall_triple_2014, (22), in the case $k=0$]. We have
@@
We input these values in $\mathcal{T}$ and use the inversion formula given in [gr, 6.576] to obtain
@@
For every $t\in\mathbb{R}$, denote $E_{t}:=E(\cdot,it)$. For every $t,s\in\mathbb{R}$ and every cusp forms $u_{k}$ and $g$, we have
@@
We start by $\langle u_{k}g,E_{t}\rangle$, which we will unfold by inputting the Fourier expansion of each cusp form and Eisenstein series. This generalizes [hei, Section 17.2], where the computation is restricted to the case $u_{k}=g$. By using the Fourier expansions (11) of $u_{k}$ and $g$, and by denoting $t_{k}$ and $t_{g}$ their respective spectral parameters, we have
@@
By executing the integration of $e(\langle\tilde{\mu}-\tilde{\nu},z\rangle)$ over $z$ first, we select only the case where $\mu=\nu$ with a weight $\mathrm{vol}(P_{\mathfrak{o}})$; indeed by [hei, p. 122] we have
@@
By executing the integration of $e(\langle\tilde{\mu}-\tilde{m},z\rangle)$ over $z$ first and using (28), we only pick up the terms where $\mu=m$ with a weight $\mathrm{vol}(P_{\mathfrak{o}})$, and this in particular removes the constant terms of the Eisenstein series’ expansion. We are therefore left with
@@
$\langle E_{t}f,E_{s}\rangle$ $=\rho_{f}(1)2^{2-(1-it)}\frac{\sqrt{|d_{K}|}}{2}\frac{|\mathcal{O}^{\times}|^{2}}{\xi_{K}(it)\xi_{K}(is)}\left(\frac{4\pi}{\sqrt{|d_{K}|}}\right)^{-it}$ (32)
$\quad\times\frac{L(\frac{it-is+1}{2},f)L(\frac{it+is+1}{2},f)}{\Gamma(it)}\prod_{\pm,\pm}\Gamma\left(\frac{it\pm(is-1)\pm it_{f}-1}{2}\right)$
@@
Assing [assing_sup-norm_2019] recently proved a stronger bound on sup-norm for Eisenstein series with exponent $3/8$, matching the quality of [young_note_2018]; however the above lemma is sufficient to our purposes and we provide an elementary proof for completeness.
@@
We follow closely [young_note_2018]. Consider the Fourier expansion given in (10) and the relation $E=\frac{|\mathcal{O}_{K}^{*}|}{2}E_{\infty}$. After applying the Cauchy-Schwarz inequality, it suffices to bound
@@
To simplify the notation, denote $L=4\pi r/\sqrt{|d_{K}|}$ and notice that both $L$ and $1/L$ are bounded in terms of $\Omega$. We need to state uniform asymptotic estimates for the $K$-Bessel function, that we encapsulate in the following lemma.
Lemma 2.9.
Let $C\gg 1$. We have
$\cosh(\pi t/2)K_{it}(u)\ll\begin{cases}t^{-1/4}(t-u)^{-1/4},\quad&\text{if }0<u<t-Ct^{1/3},\\ t^{-1/3},\quad&\text{if }|u-t|\leq Ct^{1/3},\\ u^{-1/4}(u-t)^{-1/4}\quad&\text{if }\tfrac{\pi}{2}t\geq u>t+Ct^{1/3},\\ u^{-1/4}(u-t)^{-1/4}\exp(-u+\pi t/2)\quad&\text{if }u>\tfrac{\pi}{2}t.\end{cases}$ (35)
Proof.
To obtain the second estimate, we borrow the expansions from Balogh [balogh_asymptotic_1967, Equation (8)] in the transition regime $|u-t|\leq Ct^{1/3}$, and note that the Airy function therein is uniformly bounded in a neighborhood of zero.
In the regime $0<u<t-Ct^{1/3}$, we appeal to Erdélyi [erdelyi, 7.13.2 (19) with $M=1$], noting that the exponential is asymptotically $\cosh(\pi t/2)$, giving the first claimed estimate.
In the regime $u>\tfrac{\pi}{2}t$, we appeal to Erdélyi [erdelyi, 7.13.2 (18) with $M=1$] in order to obtain extra decay. Asymptotically expanding the argument in the exponential gives that the exponential is $\ll\cosh(\pi t/2)\exp(-u+\pi t/2)$ as claimed in the fourth stated estimate.
In the regime $\tfrac{\pi}{2}t>u>t+Ct^{1/3}$, we appeal to Erdélyi [erdelyi, 7.13.2 (18) with $M=1$] . Let $x=u/t$. In the range $x\geq 1$, the argument of the exponential is $t\left(\sqrt{x^{2}-1}+\arcsin(1/x)\right)$. We apply the identity $\pi/2-\arcsin(1/x)=\arctan\left(\sqrt{x^{2}-1}\right)$, valid for $x\geq 1$, to rewrite the argument as $\pi t/2-t(v-\arctan(v))$, for $v=\sqrt{x^{2}-1}$. We recall that the contribution of the argument $\pi t/2$ exactly corresponds to the one of $1/\cosh(\pi t/2)$. Finally, we notice that $v-\arctan{v}\geq 0$, for $v\geq 0$, since the derivative of the left hand side is non-negative and the inequality holds for $v=0$. ∎
@@
using Shiu’s bound [shiu_brun-titschmarsh_1980] on the function $\tilde{\sigma}_{0}(n)=\frac{1}{\sqrt{n}}\sum_{|\omega|^{2}=n}\sigma_{0}(\omega)$. These terms then give the same bound as (36), plus an additional term that is $\ll_{\Omega}t^{-1/3+\varepsilon}$. The contribution from the remaining range with $u>2t$ is easily bounded thanks to the exponential contribution in (35). ∎
@@ 3. Proof of the theorem
We have now all the tools necessary to prove Theorem 1.2. We follows the strategy of Hua, Huang and Li [hua_joint_2024].
@@ 3.1. Spectral expansion and first truncation
Using the Cauchy-Schwarz inequality, the fact that $\|\Delta^{\ell}\psi\|_{1}\ll 1$ and $\|E_{t}\|_{\infty}\ll t^{1+\varepsilon}$ from Lemma 2.7, we obtain $\langle\psi,E_{t}\rangle\ll\lambda_{t}^{-\ell}$ for any $\ell>0$.
@@
$\langle\psi,f^{2}g\rangle$ $=\frac{\langle\psi,1\rangle\langle 1,f^{2}g\rangle}{\mathrm{vol}(\Gamma\backslash\mathbb{H}^{3})}+\delta_{t_{g}\ll t_{f}^{o(1)}}\langle\psi,g\rangle\langle 1,f^{2}g^{2}\rangle+\sum_{\begin{subarray}{c}u_{k}\neq g\\ t_{k}\ll\max(t_{f},t_{g})^{o(1)}\end{subarray}}\langle\psi,u_{k}\rangle\langle u_{k},f^{2}g\rangle$
@@ 3.2. Reduction to triple products
Many terms in (3.1) display inner products involving four forms. In order to make them amenable to the Watson-Ichino formula, we apply further the spectral expansion from Lemma 2.2 to the terms $u_{k}g$ and $E_{t}g$, which allows to reduce the expression to products of three forms at the cost of an extra spectral summation. We obtain